TeXML FAQ

June 23, 1999

What can I do with TeXML?

You can create typeset documents from XML markup.

What kind of problem might TeXML help me solve?

The power of XML is that it allows you to identify information with a domainspecific, domain-standard tags. This is what we refer to as “content markup.” The elements describe the data itself, not the presentation of the data.

Bibliographic content, for example, might have the entry, “<title>War and Peace</title>.” That is content markup. For presentation in a browser we might write “<i>War and Peace</i>.” For presentation in a journal we might write “{\it War and Peace}.”

Sooner or later you will want to produce printed or typeset pages from some XML content. TeXML enables you to do so using TeX.

Are there other applications?

This FAQ was written with XML markup and transformed to HTML markup. It is also available as PostScript generated through TeXML. The materials available here demonstrate the transform to PostScript from the same markup used to produce the HTML. One markup. Two presentations– HTML and print.

The materials also include an example of MathML markup transformed into TeXML, with PostScript generated using TeX.

What do I need to use TeXML?

You will need the following in addition to the files provided here:

• An XSL implementation. This version of TeXML was tested with Lotus XSL version 0.16.5 available from the IBM AlphaWorks web site, http://www.alphaworks.ibm.com/ • A java runtime implementation. You will find these at http://www.javasoft.com/products/jdk/1.1/jre/.

This version of TeXML was tested with JDK version 1.1.7B.

• An XML parser. TeXMLatte uses the XML4J parser available from the IBM AlphaWorks web site, http://www.alphaworks.ibm.com/. This version of TeXML was tested with XML4J version 2.0.6.

• A LaTeX implementation. You will find references to these at the TeX User’s Group, http://www.tug.org/.

What is TeX?

TeX is a language for typesetting. TeX documents are ASCII coded documents which contain typesetting commands interwoven with the content to be typeset. Implementations of TeX are widely available for most operating systems. The same TeX source may be processed on any of these systems to produce the same typeset output.